The present value of $2,000 to be paid one year later with earnings of 10% is?
Answer #1

Answer : 1,818.18 (or) 1,818.2

Present Value = 1/(1+r)^n * Payment = 2,000*(1/(1+10%)^1 = 1,818.18 (or) 1,818.2

Answers is Sensitive to the Round off Factor . Please choose the above options . According to the Requirements.